What pressure will be exerted by 25 g of CO2 at a temperature of 25 degrees celsius and a volume of 500 ml


1
Expert's answer
2022-06-07T07:48:02-0400

Mr ( CO2) = 44 g/mol

T = 25°C = 298 K

V = 500 ml = 0,5 l

m(CO2) = 25 g

P = ?


1.

n = m/M = 25 g / 44 g/mol =0,5682 mol CO2

2.

PV = nRT

R — universal gas constant = 8,31

P = nRT / V

P = 0,5682 * 8,31 * 298 / 0,5 = 2814 kPa

ANSWER: 2814 kPa